OpenStudy (anonymous):

Give an example of a ring in which the cancellation law does not hold,that is find a ring R with nonzero elements a,x,y such that ax=ay but x #y (please be sure to provide a specific example in the ring you found

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Well, what are a few examples of rings?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

I am not sure of this question

OpenStudy (anonymous):

Do you know the definition of a ring?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

I don't quite understand -- I assume you are in some abstract algebra course?

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

Do you know the definition of a ring?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so give me a few examples of rings.

OpenStudy (anonymous):

ring how has at least two elements and has all the properties without mult invers ,mult identity and mult commit vie

OpenStudy (anonymous):

I have no idea what that sentence is meant to say, I'm afraid. Just give me one example of a ring.

OpenStudy (anonymous):

and a commitive ring is ring whose multiplication is commutative

OpenStudy (anonymous):

Not definitions, examples.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I just have that R(X) AND F(X) IS RING

OpenStudy (anonymous):

I DONT HAVE EXAMPLE JUST DEFINATION

OpenStudy (anonymous):

Well they don't do much good if you don't understand what the words mean. Do you understand what a group is?

OpenStudy (anonymous):

I really understand ring what is mean but I do not understand whats the question mean

OpenStudy (anonymous):

The question is asking for an example of a ring with elements a,b, and c such that \[ a\otimes b = a\otimes c \] but \[ b \neq c \]

OpenStudy (anonymous):

And respectfully, if you can't give an example of a ring, then you don't know what a ring is.

OpenStudy (anonymous):

yes because I do not have example of that and I asked because I do not know about it just know the definition

OpenStudy (anonymous):

Well, like I asked before, do you know what a group is?

OpenStudy (anonymous):

And can you give me an example of a group?

OpenStudy (anonymous):

i think z7 is aring

OpenStudy (anonymous):

z mode 7

OpenStudy (abb0t):

wtf is this?

OpenStudy (anonymous):

Yes, it is. Z / 7 is a ring. So here's the question: Does every multiplication have an inverse?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, which ones are not invertible?

OpenStudy (anonymous):

you ask about z mode 7 or what

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

because its integer and there is no invers

OpenStudy (anonymous):

Okay. Let's say I have an integers "a", "b", and "c". Is it possible that a*b = a*c, with b not equal to c?

OpenStudy (anonymous):

you mean a times b ?

OpenStudy (anonymous):

okay i understood

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes that is possible of course

OpenStudy (anonymous):

like 4*4=8*2

OpenStudy (anonymous):

but what is the question want find example in ring then we have to prove does not equal

OpenStudy (anonymous):

Notice that I'm saying \[ a\otimes b = a\otimes c \] the integer "a" is common to both terms.

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

does have to be integer

OpenStudy (anonymous):

yes, z/7, remember?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am really confused about that

OpenStudy (anonymous):

About what?

OpenStudy (anonymous):

about how can i solve it!

OpenStudy (anonymous):

There are many very simple examples. Consider the ring of integers. \[0 \otimes 4 = 0 \otimes 7 \] right?

OpenStudy (anonymous):

but the question say with nonzero elements

OpenStudy (anonymous):

I understand, we're getting there.

OpenStudy (anonymous):

So we can't find elements like that in the set of integers. What's another simple example? How about matrices?

OpenStudy (anonymous):

iHave idea

OpenStudy (anonymous):

if we take (1)mod 7*(6)mode 7=(1)mode 7 *(13) when we 13-7=6

OpenStudy (anonymous):

do you think that's work?

OpenStudy (anonymous):

13 is not an integer mod 7, though. I encourage you to think about matrices instead.

OpenStudy (anonymous):

how with matrix

OpenStudy (anonymous):

if A and B are nonzero matrices, is it possible that \[A \otimes B = 0\] where 0 is obviously the zero matrix?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What about these matrices \[\left[\begin{matrix}1 &1 \\ 1 & 1\end{matrix}\right]\cdot \left[\begin{matrix}1& 1\\-1 & -1\end{matrix}\right]\]

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

What if I changed the second matrix to\[\left[\begin{matrix}1 & 2 \\ -1 & -2\end{matrix}\right]\]

OpenStudy (anonymous):

Would that still be zero?

OpenStudy (anonymous):

yeeeees still zero

OpenStudy (anonymous):

So we've found matrices such that \[A\otimes B = A\otimes C\] and A,B,and C are all nonzero.

OpenStudy (anonymous):

is it include which group the matrices

OpenStudy (anonymous):

I don't know what that means.

OpenStudy (anonymous):

i mean in this question say what is the number you choose have to be ring

OpenStudy (anonymous):

The set of all 2x2 matrices is a group, whose composition law is element-wise addition. When you impose the additional operation of matrix multiplication, which distributes over the element-wise addition, you form a ring.

OpenStudy (anonymous):

oh okay i understood

OpenStudy (anonymous):

thank you so much for your cooperation. i appreciate that

OpenStudy (anonymous):

No problem.